ADNAN MUHAMMED

Grade 12,

A body of 2kg collides with a weightless spring of force constant 4 N/m.The body compresses the spring by 1m from its rest position.Find the speed of the body at instant of collision.(Coefficient of friction-0.1)

A body of 2kg collides with a weightless spring of force constant 4 N/m.The body compresses the spring by 1m from its rest position.Find the speed of the body at instant of collision.(Coefficient of friction-0.1)

Grade:11

1 Answers

India VK
10 years ago

Use energy conservation principle.

 

At max. compression of the spring: 

1. PE of the spring = 0.5 * 4 * 1*1 = 2 J

2. Energy lost via friction = 2*10*0.1 = 2 J

 

So the KE of the block at the instant of collision with the spring = 2 + 2 = 4 J

 

Hence its speed = 2 m/s